In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solution phase reaction.

Since the cup itself can absorb energy, a separate experiment is needed to determine the heat capacity of the calorimeter. This is known as calibrating the calorimeter and the value determined is called the calorimeter constant.

One way to do this is to use a common metal of known heat capacity. In the laboratory a student heats 99.35 grams of copper to 98.77 °C and then drops it into a cup containing 81.42 grams of water at 24.22 °C. She measures the final temperature to be 31.71 °C.

Using the accepted value for the specific heat of copper (See the References tool), calculate the calorimeter constant.
Answer #1

m(water) = 81.42 g
T(water) = 24.22 oC
C(water) = 4.184 J/goC
m(copper) = 99.35 g
T(copper) = 98.77 oC
C(copper) = to be calculated

We will be using heat conservation equation

use:
heat lost by copper = heat gained by water
m(copper)*C(copper)*(T(copper)-T) = m(water)*C(water)*(T-T(water))
99.35*C(copper)*(98.77-31.71) = 81.42*4.184*(31.71-24.22)
6662.411*C(copper) = 2551.553
C(copper)= 0.383 J/goC
Answer: 0.383 J/goC